OpenStudy (anonymous):

what equation does not have the same solutions as 3x^2=6x-9

OpenStudy (anonymous):

\[x=\pi\]

OpenStudy (anonymous):

transcendental number

OpenStudy (anonymous):

whats a transcendental number

OpenStudy (anonymous):

so out of these choices which one is correct 1. 3x^2-6x+9=0 2.3x^2-6x=-9 3.3x62-6x-9=0 4.x^2-2x=3=0